Lock, Statistics: Unlocking the Power of Data, 2e Grace Period: 1 days left | Register Now PRINTER VERSIONBACK NEXT ASSIGNMENT RESOURCES Sum18- Assignment 3 Chapter 2, Section 6, Exercise 208acdefgh Concentration or coz in the Atmosphere Levels of carbon dioxide (CO2) in the atmosphere are rising rapidly, far above any levels ever before recorded. Levels were around 278 parts per million in 1800, before the Industrial Age, and had never, in the hundreds of thousands of years before that, gone above 300 ppm. Levels are now nearing 400 ppm. Table 1 shows the rapid rise of CO2 concentrations over the last 50 years, also available in CarbonDioxide. We can use this information to predict CO2 levels in different years. Year 1960 1965 1970 1975 1980 1985 1990 1995 2000 2005 2010 C02 316.91 320.04 325.68 331.08 338.68 345.87 354.16 360.62 369.40 379.76 389.78 Chapter P Section 1, Chapter P, Section 1, Exercise 067 Chapter P, Section 3, Table 1 Concentration of carbon dioxide in the atmosphere Review Results by Study Objective Click here for the dataset associated with this question Dr. Pieter Tans, NOAAVESRL, www.esrl.noaa.gov/gmd/ccgg/trends. Values recorded at the Mauna Loa Observatory in Hawaii. (a) What is the explanatory variable? Year O CO2 level
Answer #1

a)
Year is an explanatory variable.

b)
r=0.9931

c)The equation of regression line is
C02 = - 2571.2 + 1.4708*Year

d)
Last Option is correct.

e)
The intercept is -2571.2

It doesn't make any sense.

f)
CO2 level in 2003 = - 2571.2 + 1.4708*2003=374.8124
CO2 level in 2020 = - 2571.2 + 1.4708*2020=399.816

g)
Residual for 2010=389.78-385.108=4.672